OpenStudy (anonymous):

Is the integral of x^3*e^(-x^4) from negative infinity to infinity is divergent?

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}x^3e ^{-x ^{4}}dx\] this is what it looks like